OpenStudy (anonymous):

Solve Equation 7/q^2-q-2+1/q+1=3/q-2.

OpenStudy (anonymous):

\[\frac{ 7 }{ q^2-q-2 }+\frac{ 1 }{ q+1}=\frac{ 3 }{ q-2 }\] Thanks in advance

jhonyy9 (jhonyy9):

q^2 -q -2 =(q+1)(q-2) hope this will help you --- so make this to common denominator and continue adding fractions ok ?